OpenStudy (anonymous):

verify: (sinx+cosx)squared=1+sin(2x)

terenzreignz (terenzreignz):

Use foil first. on \[\large (\sin x + \cos x)^2\]

OpenStudy (anonymous):

can you show me how to solve it from the right side?

terenzreignz (terenzreignz):

Sure. First, how do you perform the FOIL method, in squaring a binomial?

terenzreignz (terenzreignz):

(x+y)² = x² + 2xy + y² right?

OpenStudy (anonymous):

yes, i got sin\[2^\]+cos\[2^\]+cos(x)sin(x)+sin(x)cos(x)

terenzreignz (terenzreignz):

Or, sin²x + cos²x + 2sinxcosx

OpenStudy (anonymous):

according to the pathagorean identity you can combine the sine and cosine Squared to equal 1

OpenStudy (anonymous):

so how do you get rid of the cosx?

terenzreignz (terenzreignz):

Yeah, so you have 1 + 2sinxcosx Now, what's the formula for sin(a+b) ?

OpenStudy (anonymous):

sinxcosy+cosxsiny

terenzreignz (terenzreignz):

That's right :) Now what's the formula for sin(2x) = sin(x+x) ?

OpenStudy (anonymous):

thank you, I forgot about that identity

terenzreignz (terenzreignz):

So you can do it from here?

OpenStudy (anonymous):

yeah, thanks

terenzreignz (terenzreignz):

No problem.